# MixGRPO: Unlocking Flow-based GRPO Efficiency with Mixed ODE-SDE

### v2

@@ Abstract
Although GRPO substantially enhances flow matching models in human preference alignment of image generation, methods such as DanceGRPO still exhibit inefficiency due to the necessity of sampling and optimizing over all denoising steps specified by the Markov Decision Process (MDP). In this paper, we propose MixGRPO, a novel framework that leverages the flexibility of mixed sampling strategies through the integration of stochastic differential equations (SDE) and ordinary differential equations (ODE). This streamlines the optimization process within the MDP to improve efficiency and boost performance. Specifically, MixGRPO introduces a sliding window mechanism, using SDE sampling and GRPO-guided optimization only within the window, while applying ODE sampling outside. This design confines sampling randomness to the time-steps within the window, thereby reducing the optimization overhead, and allowing for more focused gradient updates to accelerate convergence. Additionally, as time-steps beyond the sliding window are not involved in optimization, higher-order solvers are supported for sampling. So we present a faster variant, termed MixGRPO-Flash, which further improves training efficiency while achieving comparable performance. MixGRPO exhibits substantial gains across multiple dimensions of human preference alignment, outperforming DanceGRPO in both effectiveness and efficiency, with nearly 50% lower training time. Notably, MixGRPO-Flash further reduces training time by 71%. Codes and models are available at MixGRPO.
@@ 1 Introduction
Recent advances (Liu et al., 2025a; Xue et al., 2025; Liu et al., 2025b; Li et al., 2025; Xu et al., 2023) in Text-to-Image (T2I) tasks have demonstrated that probability flow models can achieve improved performance by incorporating Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) strategies during the post-training stage to maximize rewards. Specifically, methods (Liu et al., 2025a; Xue et al., 2025) based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024), have recently been studied, achieving optimal alignment with human preferences.
Current GRPO methods in probability flow models, e.g., Flow-GRPO (Liu et al., 2025a), DanceGRPO (Xue et al., 2025), leverage Stochastic Differential Equations (SDE) sampling at every denoising step to introduce randomness, addressing reliance on stochastic exploration in RLHF. They model the entire denoising process as a Markov Decision Process (MDP) in a stochastic environment, using GRPO to optimize the state-action sequence. However, the need to optimize all denoising steps not only increases overhead but also leads to inconsistent gradient descent, resulting in inefficient training. Specifically, to compute the policy ratio, it is essential to perform full-step sampling independently with the old policy model $\pi_{\theta_{\text{old}}}$ and new one $\pi_{\theta}$. While DanceGRPO proposes to randomly select a subset of steps at a fixed ratio to optimize, our empirical analysis in Figure 1 demonstrates a substantial performance degradation as the subset size is reduced.
Figure 1: Performance comparison for different numbers of denoising steps optimized. The performance improvement of DanceGRPO relies on more steps optimized. MixGRPO achieves optimal performance while requiring only 4 steps.
To address these issues, we propose MixGRPO, which achieves more efficient stochastic exploration while enabling optimization over fewer denoising steps. Specifically, we employ a mixed ODE-SDE strategy, applying SDE sampling to a denoising sub-interval and Ordinary Differential Equations (ODE) sampling to the rest, confining randomness to the SDE interval. In this way, fewer time-steps are needed for GRPO optimization, without compromising image quality for reward computation. Besides, we introduce a sliding window strategy for the SDE interval that moves along the denoising steps as the training process progresses. Unlike random selection, this scheduling strategy orders optimization from high to low denoising levels, which aligns with the intuition of applying temporal discount factors to rewards in Reinforcement Learning (RL) (Pitis, 2019; Amit et al., 2020; Hu et al., 2022b). MixGRPO prioritizes optimizing the initial timesteps, which involve the most significant noise removal and entail a larger exploration space (see Figure 2). As training progresses, MixGRPO continuously moves the SDE window to narrow the exploration space and achieve global optimization. Finally, we find that higher-order ODE solvers, e.g., DPMSolver++ (Lu et al., 2022b), can significantly accelerate training-time sampling, with negligible performance degradation, as there is no need for the posterior probability distribution after the sliding window.
@@
We trained and evaluated MixGRPO by using HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), ImageReward (Xu et al., 2023), and Unified Reward (Wang et al., 2025) as reward models (RMs) and metrics. We also quantified the overhead in terms of the number of function evaluations (NFE) and time consumption during training. During the training process, we fine-tuned based on FLUX.1-dev (Labs, 2024) and compared performance using either a single RM or multiple RMs as guidance, assessing the results for both in-domain and out-of-domain metrics. Specifically, trained and evaluated on the HPDv2 dataset (Wu et al., 2023), MixGRPO outperforms DanceGRPO across all metrics, particularly improving the ImageReward (Xu et al., 2023) from 1.088 to 1.629, surpassing DanceGRPO’s score of 1.436, while generating images with enhanced semantic quality, aesthetics, and reduced distortion. Furthermore, MixGRPO reduces the training time of DanceGRPO by nearly 50%. In addition, MixGRPO-Flash utilizes DPMSolver++ (Lu et al., 2022b) to accelerate the sampling of $\pi_{\theta_{\text{old}}}$, reducing training time by 71%.
To summarize, the key contributions of our work are outlined below:
@@
  • We propose a mixed ODE-SDE GRPO training framework for flow-based models, which alleviates the overhead bottleneck by streamlining the optimization process within the MDP.
  • We introduce a sliding window strategy to sample the denoising timesteps for model optimization, aligning with the RL intuition of transitioning from harder to easier search spaces, significantly enhancing performance.
@@
  • Our method enables the use of higher-order ODE solvers to accelerate $\pi_{\theta_{\text{old}}}$ sampling during GRPO training, achieving more significant speed improvements with comparable performance.
@@
  • Comprehensive experiments were carried out on multiple rewards and the results demonstrate that MixGRPO achieves substantial gains on various evaluation metrics, while significantly reducing training overhead.
Figure 2: Visualization of t-SNE (Van der Maaten & Hinton, 2008) for images sampled with different strategies. Employing SDE sampling in the early stages of the denoising process results in a more discrete data distribution.
@@ 2.1 RL For Image Generation
Inspired by Proximal Policy Optimization (PPO) (Schulman et al., 2017), early works (Fan & Lee, 2023; Black et al., 2023; Fan et al., 2023; Lee et al., 2023) integrated reinforcement learning (RL) into diffusion models by optimizing the score function (Song et al., 2020b) through policy gradient methods, thereby enabling the generation of images that better align with human preferences. Subsequently,  Wallace et al. (2024) introduced offline-Direct Preference Optimization (DPO) to T2I tasks for the first time. This allows diffusion models to directly learn from human feedback and validates its effectiveness on large-scale models. Due to the tendency of offline win-lose pair data to shift the model away from its original distribution, some works (Yuan et al., 2024; Liang et al., 2025) have adopted online methods, continuously adjusting sampling trajectories through step-aware preference models during training to achieve improved performance. Recently, GRPO-based works e.g.,  Tong et al. (2025), Flow-GRPO (Liu et al., 2025a) and DanceGRPO (Xue et al., 2025), have elevated RL-enhanced image generation to new heights. Specifically, Flow-GRPO (Liu et al., 2025a) and DanceGRPO (Xue et al., 2025) introduced GRPO to flow matching models, allowing divergent sampling by transforming the ODE into an equivalent SDE. They also identified the overhead caused by full-step sampling within a group as a bottleneck and sought to address it by reducing or randomly selecting denoising steps. However, these approaches do not fundamentally address the issue. We hope to delve into the essence of GRPO on the probability flow and provide deeper insights through mixed sampling techniques and optimization scheduling.
@@ 2.2 Sampling Methods for Probability Flow
DDPM (Ho et al., 2020) first proposed training a probabilistic model to reverse each step of noise corruption and utilized probability flow SDE for sampling, enabling the generation of realistic images. However, this often requires thousands of steps, resulting in significant overhead. DDIM (Song et al., 2020a) introduced deterministic sampling and proposed a probability ODE sampling approach, reducing the number of sampling steps to around 100. Subsequently, inspired by the Fokker-Planck equation (Risken & Risken, 1996), Song et al. (2020b) established a unification of SDE and ODE sampling methods from the perspective of the score function. Then, more higher-order ODE solvers were proposed e.g., DPM-Solver (Lu et al., 2022a) and DPM-Solver++ (Lu et al., 2022b), which utilize multistep methods for differential discretization. These approaches significantly reduce the number of sampling steps to around 10 while preserving accuracy. Higher-performance solvers (Zheng et al., 2023; Zhao et al., 2023) continue to be proposed; however, the gains are relatively marginal and have ultimately been replaced by the distillation method (Salimans & Ho, 2022; Yin et al., 2024). During the same period, flow matching models (Lipman et al., 2022; Esser et al., 2024) simplified and stabilized training by predicting the vector field velocity, enabling deterministic sampling with ODEs under 50 steps. Recent theoretical works (Gao et al., 2024; Albergo et al., 2023) has proven that the sampling method of flow matching is equivalent to DDIM, and demonstrated that flow matching models share the same equivalent SDE and ODE formulations as diffusion models. This provides important theoretical support and insights for our work, and we may explore interleaved sampling of SDE and ODE in probability flow models as a potential approach.
@@ 3.1 Mixed ODE-SDE Sampling in GRPO
According to Flow-GRPO (Liu et al., 2025a), the SDE sampling in flow matching can be framed as a Markov Decision Process (MDP) $\left(\mathcal{S},\mathcal{A},\rho_{0},P,\mathcal{R}\right)$ in a stochastic environment. The agent produces a trajectory during the discrete sampling process defined as $\Gamma=(\mathbf{s}_{0},\mathbf{a}_{0},\mathbf{s}_{1},\mathbf{a}_{1},\ldots,\mathbf{s}_{T},\mathbf{a}_{T})$, where the reward is provided only at the final step by the reward model, specifically $\mathcal{R}(\mathbf{s}_{i},\mathbf{a}_{i})\overset{\triangle}{=}R(\mathbf{x}_{T},c)$ if $i=T$, and 0 otherwise.
@@
In MixGRPO, we propose a hybrid sampling method that combines SDE and ODE. MixGRPO defines a time interval $S=[t_{l},t_{r})\in[0,1)$, which corresponds to a subinterval of denoising timesteps, such that $0\leq l<r\leq T$ and $t_{i}=\frac{i}{T}$. We use SDE sampling within the interval $S$ and ODE sampling outside, while $S$ shifts along the denoising direction throughout the training process (See Figure 2). MixGRPO restricts the agent’s stochastic exploration space to the interval $S$, shortening the sequence length of the MDP to a subset $\Gamma_{\text{MixGRPO{}}}=(\mathbf{s}_{l},\mathbf{a}_{l},\mathbf{s}_{l+1},\mathbf{a}_{l+1},\ldots,\mathbf{s}_{r},\mathbf{a}_{r})$ and requires reinforcement learning (RL) optimization only on this subset:
@@
$\max_{\theta}\mathbb{E}_{\Gamma_{\text{MixGRPO}}\sim\pi_{\theta}}\left[\sum_{t\in S}\big(\mathcal{R}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)-\beta D_{KL}\left(\pi\left(\cdot|\mathbf{s}_{t}\right)||\pi_{\text{ref}}\left(\cdot|\mathbf{s}_{t}\right)\right)\big)\right],$ (1)
@@
where $\mathcal{R}(\mathbf{s}_{i},\mathbf{a}_{i})\overset{\triangle}{=}R(\mathbf{x}_{T},c)$ if $i=r$, and 0 otherwise. MixGRPO reduces computational overhead while also lowering the difficulty of optimization. Next, we derive the specific sampling form and optimization objective of MixGRPO.
For a deterministic reverse probability flow ODE (Song et al., 2020b), it takes the following form:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}),$ (2)
@@
where $q_{t}(\mathbf{x}_{t})$ represents the evolution process of the reverse probability distribution from 0 to $T$. $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}$ is the score function at time $t$. According to the Fokker-Planck equation (Risken & Risken, 1996; Øksendal, 2003),  Song et al. (2020b) has demonstrated that Eq. (2) has the following equivalent probability flow SDE, which maintains the same marginal distribution at each time $t$:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}+g(t)\frac{\mathrm{d}\mathbf{w}}{\mathrm{d}t},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}).$ (3)
@@
In MixGRPO, we mix ODE and SDE for sampling, which has the same convergence as using only ODE sampling (a detailed proof in Appendix A). The specific form is as follows:
@@
In particular, for Flow Matching (FM) (Lipman et al., 2022), especially the Rectified Flow (RF) (Liu et al., 2022), the sampling process can be viewed as a deterministic ODE:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\mathbf{v}_{t}.$ (5)
Eq. (5) is actually a special case of the Eq. (2) with $\mathbf{v}_{t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}$. So we can derive the ODE-SDE hybrid sampling form for RF as follows:

### v3

@@ Abstract
Although GRPO substantially enhances flow matching models in human preference alignment of image generation, methods such as FlowGRPO and DanceGRPO still exhibit inefficiency due to the necessity of sampling and optimizing over all denoising steps specified by the Markov Decision Process (MDP). In this paper, we propose MixGRPO, a novel framework that leverages the flexibility of mixed sampling strategies through the integration of stochastic differential equations (SDE) and ordinary differential equations (ODE). This streamlines the optimization process within the MDP to improve efficiency and boost performance. Specifically, MixGRPO introduces a sliding window mechanism, using SDE sampling and GRPO-guided optimization only within the window, while applying ODE sampling outside. This design confines sampling randomness to the time-steps within the window, thereby reducing the optimization overhead, and allowing for more focused gradient updates to accelerate convergence. Additionally, as time-steps beyond the sliding window are not involved in optimization, higher-order solvers are supported for faster sampling. So we present a faster variant, termed MixGRPO-Flash, which further improves training efficiency while achieving comparable performance. MixGRPO exhibits substantial gains across multiple dimensions of human preference alignment, outperforming DanceGRPO in both effectiveness and efficiency, with nearly 50% lower training time. Notably, MixGRPO-Flash further reduces training time by 71%. [^1] Code is available in https://github.com/Tencent-Hunyuan/MixGRPO.
@@ 1 Introduction
Recent advances [17, 42, 18, 41, 48] in Text-to-Image (T2I) tasks have demonstrated that probability flow models can achieve improved performance by incorporating Reinforcement Learning from Human Feedback (RLHF) [25] strategies during the post-training stage to maximize rewards. Specifically, methods [17, 42] based on Group Relative Policy Optimization (GRPO) [32], have recently been studied, achieving optimal alignment with human preferences.
Current GRPO methods in probability flow models, such as Flow-GRPO and DanceGRPO, combine a rollout phase followed by an online optimization phase. During rollouts, they employ Stochastic Differential Equation (SDE) sampling at each denoising step to introduce randomness, thereby enabling the stochastic exploration required by RLHF. This allows the entire denoising process to be framed as a Markov Decision Process (MDP) in a stochastic environment, where GRPO is then applied to optimize the complete state-action sequence in the online optimization phase. However, the requirement to optimize the entire denoising steps introduces two challenges, prohibitive computational overhead and unfocused and inefficient optimization. i) The overhead arises because computing the policy ratio requires separate full-trajectory sampling from both the old ($\pi_{\theta_{\text{old}}}$) and new policies ($\pi_{\theta}$). Although DanceGRPO proposes a workaround by optimizing a random subset of steps, our findings in Figure 1 reveal a critical trade-off: reducing the subset size to save computation severely compromises model performance. ii) Early-step gradients prioritize global structure while late-step gradients focus on fine details, leading to conflicting update signals. This may result in an inefficient training process.
Figure 1: Comparison of our MixGRPO and DanceGRPO with varying denoising steps to be optimized. MixGRPO achieves higher performance with lower overhead.
To address these issues, we propose MixGRPO, a method that optimizes a strategic subset of denoising steps to drastically reduce computational overhead while ensuring a more focused and efficient optimization. Specifically, we employ a mixed ODE-SDE strategy, applying SDE sampling to a denoising sub-interval and Ordinary Differential Equations (ODE) sampling to the rest.
Figure 2: (Left) Methods Comparison. By employing the mixed ODE-SDE and sliding window strategy, MixGRPO reduces the number of optimized timesteps and achieves more efficient training than DanceGRPO [42]. (Right) SDE Exploration Analysis. As training progresses, MixGRPO gradually slides the SDE window from a high-noise to a low-noise regime, causing the exploration space to shrink and the distribution of sampled images to transition from dispersed to concentrated.
The policy optimization is confined only to the stochastic SDE sub-interval. This interval operates as a sliding window, systematically advancing from low-SNR (signal-to-noise ratio) steps to high-SNR denoising steps as training progresses. This curriculum-based approach is conceptually analogous to temporal discounting in RL [26, 2, 10], as it prioritizes the optimization of initial, high-impact steps that establish global structure from a vast exploration space (Figure 2 Right), before progressively shifting focus to the refinement of local details. This selective optimization not only reduces the computational burden but also fosters a more dedicated exploration. Furthermore, this decoupling allows us to accelerate the deterministic ODE portions with fast solvers (e.g., DPMSolver++ [22]), since their posterior distributions are irrelevant to the optimization, thereby reducing rollout time without compromising final image quality.
@@
To evaluate the performance and efficiency of MixGRPO, we conduct comprehensive experiments by implementing our method within the frameworks of both Flow-GRPO [17] and DanceGRPO [42]. We use a set of reward models (HPS-v2.1 [40], Pick Score [12], ImageReward [41], Unified Reward [37], etc) that serve as both training guidance and evaluation metrics. The evaluation compares single versus multi-RM guidance and assesses both in-domain and out-of-domain performance. Alongside performance, computational overhead is tracked via the number of function evaluations (NFE) and training time. Specifically, trained and evaluated on the HPDv2 dataset [40], MixGRPO outperforms DanceGRPO across all metrics, particularly improving the ImageReward [41] from 1.088 to 1.629, surpassing DanceGRPO’s score of 1.436, while generating images with enhanced semantic quality, aesthetics, and reduced distortion. Furthermore, MixGRPO reduces the training time of DanceGRPO by nearly 50%. In addition, MixGRPO-Flash utilizes DPMSolver++ [22] to accelerate the sampling of $\pi_{\theta_{\text{old}}}$, reducing training time by 71%.
To summarize, the key contributions are outlined below:
@@
  • We propose a mixed ODE-SDE framework for GRPO training with flow-based models. This approach alleviates computational overhead by confining the stochastic optimization to a specific sub-interval of the Markov Decision Process.
  • We introduce an SDE sliding window to dynamically schedule the optimization steps. By guiding the optimization from broad exploration to fine-grained refinement, this strategy significantly enhances performance.
@@
  • Our hybrid framework enables the use of high-order ODE solvers to accelerate the deterministic sampling of $\pi_{\theta_{\text{old}}}$ during GRPO training. This yields substantial speedups with negligible performance degradation.
@@
  • Comprehensive experiments across multiple frameworks demonstrate the versatility of MixGRPO. Our method consistently achieves substantial performance gains while significantly reducing training overhead, highlighting its broad applicability and effectiveness.
@@ 2.1 RL For Image Generation
Inspired by Proximal Policy Optimization (PPO) [31], early works [5, 3, 6, 14] integrated reinforcement learning (RL) into diffusion models by optimizing the score function [34] through policy gradient methods, thereby enabling the generation of images that better align with human preferences. Subsequently,  [36] introduced offline-Direct Preference Optimization (DPO) to T2I tasks for the first time. This allows diffusion models to directly learn from human feedback and validates its effectiveness on large-scale models. Due to the tendency of offline win-lose pair data to shift the model away from its original distribution, some works [45, 15] have adopted online methods, continuously adjusting sampling trajectories through step-aware preference models during training to achieve improved performance. Recently, GRPO-based works e.g.,  [35], Flow-GRPO [17] and DanceGRPO [42], have elevated RL-enhanced image generation to new heights. Specifically, they introduced GRPO to flow matching models, allowing divergent sampling by transforming the ODE into an equivalent SDE. DanceGRPO [42] also identified the overhead caused by full-step optimization as a bottleneck and sought to address it by reducing or randomly selecting a subset. However, this approach does not fundamentally address the issue. We hope to delve into the essence of GRPO on the probability flow and provide deeper insights through mixed sampling techniques and optimization scheduling.
@@ 2.2 Sampling Methods for Probability Flow
The development of generative sampling began with DDPM [8], which used a slow, SDE-based probability flow requiring thousands of steps. To accelerate this, DDIM [33] introduced a deterministic ODE-based approach, reducing sampling steps to around 100. This SDE / ODE duality was later unified by score-based models [34], paving the way for higher-order ODE solvers like DPM-Solver [21, 22] that reduced steps to 10. Higher-performance solvers [47, 46] continue to be proposed; however, the gains are relatively marginal and have ultimately been replaced by the distillation method [30, 44]. Concurrently, flow matching models [16, 4] simplified training by directly predicting the vector field, also enabling fast ODE sampling. Crucially, recent theoretical work [7, 1] has proven that flow matching and diffusion models share equivalent SDE and ODE formulations. This unification provides the theoretical foundation for our work, in which we explore an interleaved SDE and ODE sampling strategy within these probability flow models, thereby confining stochasticity to the SDE intervals in RL optimization, which shortens the effective MDP horizon and reduces training overhead.
@@ 3.1 Mixed ODE-SDE Sampling in GRPO
According to Flow-GRPO [17], the SDE sampling in flow matching can be framed as a Markov Decision Process (MDP) $\left(\mathcal{S},\mathcal{A},\rho_{0},P,\mathcal{R}\right)$. The agent produces a trajectory during the discrete sampling process defined as $\Gamma=(\mathbf{s}_{0},\mathbf{a}_{0},\mathbf{s}_{1},\mathbf{a}_{1},\ldots,\mathbf{s}_{T},\mathbf{a}_{T})$, where the reward is provided only at the final step by the reward model, specifically $\mathcal{R}(\mathbf{s}_{i},\mathbf{a}_{i})\triangleq R(\mathbf{x}_{T},c)$ if $i=T$, and 0 otherwise.
@@
In MixGRPO, we propose a sampling method that combines SDE and ODE. MixGRPO defines a time interval $S=[t_{l},t_{r})\subseteq[0,1)$, which corresponds to a subset of denoising timesteps, such that $0\leq l<r\leq T$ and $t_{i}=\frac{i}{T}$. We use SDE sampling within the interval $S$ and ODE sampling outside, while $S$ shifts along the denoising direction throughout the training process (See Figure 2 Left). MixGRPO restricts the stochastic exploration to the interval $S$, shortening the sequence length of the MDP to a subset $\Gamma_{\text{MixGRPO{}}}=(\mathbf{s}_{l},\mathbf{a}_{l},\mathbf{s}_{l+1},\mathbf{a}_{l+1},\ldots,\mathbf{s}_{r},\mathbf{a}_{r})$ and requires reinforcement learning (RL) optimization only within $S$:
@@
$\max_{\theta}\mathbb{E}_{\Gamma_{\text{MixGRPO}}\sim\pi_{\theta}}\bigg[$ $\sum_{t\in S}\bigg(\mathcal{R}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)$ (1)
$-\beta D_{\text{KL}}\Big(\pi\left(\cdot|\mathbf{s}_{t}\right)||\pi_{\text{ref}}\left(\cdot|\mathbf{s}_{t}\right)\Big)\bigg)\bigg],$
@@
MixGRPO reduces training overhead while also enabling more concentrated gradient updates. Next, we derive the specific sampling form and optimization objective.
For a deterministic reverse probability flow ODE [34], it takes the following form:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})},$ (2)
@@
where $q_{t}(\mathbf{x}_{t})$ represents the evolution process of the reverse probability distribution from 0 to $T$. $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}$ is the score function at time $t$. According to the Fokker-Planck equation [28, 24],  [34] has demonstrated that Eq. (2) has the following equivalent probability flow SDE, which maintains the same marginal distribution at each time $t$:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}+g(t)\frac{\mathrm{d}\mathbf{w}}{\mathrm{d}t}.$ (3)
@@
In MixGRPO, we mix ODE and SDE for sampling, which has the same convergence as using only ODE sampling (a detailed proof in Supplementary). The specific form is as follows:
$\mathrm{d}\mathbf{x}_{t}\!=\!\begin{cases}\left[f(\mathbf{x}_{t},t)\!-g^{2}(t)\mathbf{s}_{t}(\mathbf{x}_{t})\right]\mathrm{d}t+g(t)\mathrm{d}\mathbf{w},\!\!&\text{if}~t\in S,\\ \left[f(\mathbf{x}_{t},t)\!-\frac{1}{2}g^{2}(t)\mathbf{s}_{t}(\mathbf{x}_{t})\right]\mathrm{d}t,\!\!&\text{otherwise},\end{cases}$ (4)
where we define the score function as $\mathbf{s}_{t}\ \triangleq\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}$. In particular, for Flow Matching (FM) [16], especially the Rectified Flow (RF) [19], the sampling process can be viewed as a deterministic ODE:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\mathbf{v}_{t}.$ (5)
Eq. (5) is actually a special case of the Eq. (2) with $\mathbf{v}_{t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\mathbf{s}_{t}(\mathbf{x}_{t})$. So we can derive mixed ODE-SDE sampling form for RF as follows:
@@
In the RF framework, the model is used to predict the velocity field, represented as $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)=\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$. Following [17], the score function is represented as $\mathbf{s}_{t}(\mathbf{x}_{t})=-\frac{\mathbf{x}_{t}}{t}-\frac{1-t}{t}\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$. The $g(t)$ is represented as the standard deviation of the noise $g(t)=\sigma_{t}$. According to the definition of the standard Wiener process, we use $\mathrm{d}\mathbf{w}=\sqrt{\mathrm{d}t}\epsilon$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. Applying Euler-Maruyama discretization for SDE and Euler discretization for ODE, we build the final denoising process in MixGRPO:
$\mathbf{x}_{t+\Delta t}=\begin{cases}\mathbf{x}_{t}+\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t)\Delta t+\sigma_{t}\sqrt{\Delta t}\epsilon,&\text{if }t\in S,\\ \mathbf{x}_{t}+\mathbf{v}_{\theta}(\mathbf{x}_{t},t)\Delta t,&\text{otherwise},\end{cases}$ (7)
where the SDE drift term for the sampling process is defined as $\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t)\triangleq\mathbf{v}_{\theta}(\mathbf{x}_{t},t)+\frac{{\sigma^{2}_{t}}\left(\mathbf{x}_{t}+\left(1-t\right)\mathbf{v}_{\theta}(\mathbf{x}_{t},t)\right)}{2t}$. According to Eq. (1), we only need to optimize GRPO [32] at the time within the interval $S$ for $N$ samples in the group. For an unbiased advantage estimation in GRPO, we constrain the source of variation across images generated from a single prompt. Specifically, we fix the position of the SDE window, thereby sourcing all stochasticity from an identical timestep range in a GRPO group. This not only concentrates the gradients within this window but is also essential for preserving the unbiased nature of the optimization. The final training objective is represented as follows:
$\mathcal{J}_{\text{MixGRPO}}(\theta)$ $=\mathbb{E}_{c\sim\mathcal{C},~\{\mathbf{x}^{i}_{T}\}^{N}_{i=0}\sim\pi_{\theta_{\text{old}}}(\cdot|c)}\left[\frac{1}{N}\sum^{N}_{i=1}\frac{1}{|S|}\sum_{t\in S}\right$. (8)
$\qquad\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\left.\text{min}\bigg(r^{i}_{t}(\theta)A^{i},\text{clip}\Big(r^{i}_{t}(\theta),1\!-\!\varepsilon,1\!+\!\varepsilon\Big)A^{i}\bigg)-\beta\mathcal{J}_{\text{KL}}\right],$
where $\varepsilon$ is the clipping hyperparameter, $r^{i}_{t}(\theta)$ is the policy ratio, $A^{i}$ is the advantage score and $\mathcal{J}_{KL}$ is the KL-regularization, defined as follows:
$r^{i}_{t}(\theta)$ $=\frac{q_{\theta}(\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},c)}{q_{\theta_{\text{old}}}({\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},c})},$ (9)
$A^{i}$ $=\frac{R\left(\mathbf{x}^{i}_{T},c\right)-\text{mean}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)}{\text{std}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)},$